5) Suppose s = 0.15, Y = 4200, K = 6100, n = 0.03, g=0.03 and δ= 0.10. This makes national saving smaller than steady-state investment, so that the amount of capital per effective worker will be falling.

7) In the graph of the Solow growth model, at any point to the left of the steady-state intersection we have national saving per effective labour greater than steady-state investment per person, causing (K/AL) to increase.

8) In the Solow growth model, an increase in the marginal propensity to consume shifts the steady-state investment line downward with the implied change in the capital stock resulting in a higher standard of living in the long run.



1
Expert's answer
2021-08-03T11:49:53-0400

5)

Suppose s = 0.15, Y = 4200, K = 6100, n = 0.03, g=0.03 and δ= 0.10. This makes national saving smaller than steady-state investment, so that the amount of capital per effective worker will be falling: Incorrect

s= 0.15

Y= 4200

K= 6100

n= 0.03

g= 0.03

δ= 0.10

"K=[\frac{s}{n+δ}]^2\\K=[\frac{0.15}{0.03−0.10}]^2\\K=[\frac{0.015}{−0.07}]^2\\K=[\frac{0.0225}{0.0049}]\\K=4.59"

With the given information, K is positive states amount of capital per effective worker will be increasing.


7)

In the graph of the Solow growth model, at any point to the left of the steady-state intersection we have national saving per effective labor greater than steady-state investment per person, causing (K/AL) to increase: Incorrect

The Solow model shows the changes in the level of output in an economy over a period of time. Solow model doesn't reflect national saving per effective labor.


8)

In the Solow growth model, an increase in the marginal propensity to consume shifts the steady-state investment line downward with the implied change in the capital stock resulting in a higher standard of living in the long run: Correct

Marginal propensity to consume (MPC) is that consumption which a consumer consumes when his income increases, as opposed to saving it. Higher the income of an individual then lower will be the MPC. As with the increased income the needs and wants of the individual are satisfied. As a result, they save more instead. At low-income levels, the MPC tends to be very high. MPC is high in low income because most or all of a person's income must be devoted to subsistence consumption.